# Averaging of quantum channels via channel-state duality

### v1

@@ 2 Averaging quantum channels over arbitrary unitary transformations
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\left(\mathcal{M}^{\Phi}_{U}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
@@
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\rho_{\operatorname{ME}}\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{T}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
@@
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
@@
is a finite dimensional regular representation of the unitary group on a representation space $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$. This is because: (i) the contragradient representation [7] $\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)={\pi}_{\operatorname{U}}^{\operatorname{in}}(U^{-1})^{T}$ is a regular representation ([8], Sec. 1.5.2), and (ii) tensor product of two regular representations is a regular representation ([8], ibid). Therefore one can apply the generalized duality theorem for group representations ([8], Sec. 4.2), which boils down to a conclusion, that the action of $\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}$, where $\pi_{\mathcal{C}}$ is a commutant of $\pi_{\operatorname{U}}$ in the algebra of endomorphisms of $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, decomposes uniquely into multiplicity-free direct sum of tensor products of irreducible representations $\xi_{\operatorname{U}}$ of the unitary group $\operatorname{U}(d)$ and irreducible representations $\xi_{\mathcal{C}}$ of a group which is represented by the commutant $\pi_{\mathcal{C}}$:
@@
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}$ (12)
@@
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}=U^{\otimes t_{\operatorname{in}}}\otimes\bar{U}^{\otimes t_{\operatorname{out}}}.$ (14)
The commutant of this representation is known to be the so-called walled Brauer algebra [7, 13, 14], the representations of which are difficult to construct [14]. Although there exist implementations of Quantum Schur Transform for this case (so called Mixed Quantum Schur Transform [15]), one can wish to avoid these representations and utilize simpler operator representations of permutations. In order to do so, let us reexpress the averaging operations in a way which removes the complex conjugation. We start from the following lemma:
@@
$J_{\mathcal{M}^{\Phi}_{U}}=\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)J^{\Gamma}_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}\right]^{\Gamma}.$ (17)
@@
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)^{\dagger}\right]^{\Gamma}$
@@
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
$=$ $\left[\mathcal{M}_{U}\left(J_{\Phi}^{\Gamma}\right)\right]^{\Gamma}=\sum_{k}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)^{\Gamma},$
@@ 3 Averaging quantum channels over non-unitary operations defined via Cartan decomposition
Let us now try to generalize averaging of quantum channels from the unitary averaging to averaging over non-unitary operations described by a non-compact symmetry group $\mathcal{G}$. Such an averaging can be defined via iterated integral over the Cartan decomposition of $\mathcal{G}$, which boils down to a representation of an arbitrary element of $\mathcal{G}$ as a composition of operations from the maximally compact component $\mathcal{K}$ of the group intertwined by an element from a maximal Abelian subgroup $\mathcal{A}$: $G=KAK^{\prime},\,K,K^{\prime}\in\mathcal{K},\,A\in\mathcal{A}$. Such idea of non-compact averaging has been introduced in [18, 5] in the context of averaging quantum states over SLOCC-type operations, and generalized to averaging over arbitrary finite-dimensional matrix representations of reductive groups in the recent work [6]. Following [6], let us take arbitrary matrix representation $\pi_{\mathcal{G}}$ of the group $\mathcal{G}$ and define the averaging of a quantum state with respect to the chosen representation as follows:
@@
in which the integration is defined over the Cartan “KAK” decomposition of $\mathcal{G}$. The integration over the two copies of maximally compact components $\mathcal{K},\mathcal{K}^{\prime}$ is performed uniformly, which means that $\int_{\mathcal{K}}\operatorname{d}\!K$ represents Haar integral over $\mathcal{K}$. Integration over $\mathcal{A}$ is performed with respect to any finite normalized measure on this (in general non-compact) manifold, and there appears one crucial assumption: elements of $\mathcal{A}$ enter the integral as normalized matrices: $AA_{\textrm{n}}$=A/||A||, in order to guarantee trace nonincreasing character of the map (20).
@@
$\mathcal{M}^{\Phi}_{G}(\rho)=\int\operatorname{d}\!G\,\mathcal{T}_{G}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{G}^{\operatorname{in}\dagger}(\rho),$ (22)
@@
Now all of the derivations (2)–(2) concerning the unitary case hold, and we have the following expression for the Choi matrix of the channel (22):
@@
In full analogy with the previous case one may want to avoid using contragradient representations in the generalized Schur-Weyl duality. Utilizing Lemma 2 the formula (24) becomes:
@@
Now the restricted representation $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$ reads in this case simply $U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}$, for $U\in\textrm{SU}(d)$, and we have, following (32):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)^{\Gamma},$ (35)
where now $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]$ are ordinary permutation operators corresponding to standard Schur-Weyl duality for $\textrm{U}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$ and the permutation group $\textrm{S}_{(t_{\operatorname{in}}+t_{\operatorname{out}})}$.

### v2

@@ 2 Averaging quantum channels over arbitrary unitary transformations
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\left(\mathcal{M}^{\Phi}_{\operatorname{U}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\rho_{\operatorname{ME}}\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{T}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
@@
is a finite dimensional rational representation of the unitary group on a representation space $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$. This is because: (i) the contragradient representation [17] $\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)={\pi}_{\operatorname{U}}^{\operatorname{in}}(U^{-1})^{T}$ is a rational representation ([29], Sec. 1.5.2), and (ii) tensor product of two rational representations is a rational representation ([29], ibid). Therefore one can apply the generalized duality theorem for group representations ([29], Sec. 4.2.1) [^2] Note that the assumption of local regularity in the general duality theorem is in the finite dimensional case equivalent to rationality of the representation, see Sec. 1.5.1 of [29]., which boils down to a conclusion, that the action of $\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}$, where $\pi_{\mathcal{C}}$ is a commutant of $\pi_{\operatorname{U}}$ in the algebra of endomorphisms of $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, decomposes uniquely into multiplicity-free direct sum of tensor products of irreducible representations $\xi_{\operatorname{U}}$ of the unitary group $\operatorname{U}(d)$ and irreducible representations $\xi_{\mathcal{C}}$ of a group which is represented by the commutant $\pi_{\mathcal{C}}$:
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}$ (14)
@@
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}=U^{\otimes t_{\operatorname{out}}}\otimes\bar{U}^{\otimes t_{\operatorname{in}}}.$ (16)
The commutant of this representation is known to be the so-called walled Brauer algebra [17, 19, 21], the representations of which are difficult to construct [21]. Although there exist implementations of Quantum Schur Transform for this case (so called Mixed Quantum Schur Transform [20, 22, 24]), one can wish to avoid these representations and utilize simpler operator representations of permutations. In order to do so, let us reexpress the averaging operations in a way which removes the complex conjugation. We start from the following lemma:
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}=\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)J^{\Gamma}_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}\right]^{\Gamma}.$ (19)
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)^{\dagger}\right]^{\Gamma}$
@@
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
$=$ $\sum_{k}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)^{\Gamma},$
@@ 3 Averaging quantum channels over non-unitary operations defined via Cartan decomposition
Let us now try to generalize averaging of quantum channels from the unitary averaging to averaging over finite-dimensional non-unitary representations of a non-compact symmetry group $\mathcal{G}$. Such an averaging can be defined via iterated integral over the Cartan decomposition of $\mathcal{G}$, which boils down to a representation of an arbitrary element of $\mathcal{G}$ as a composition of operations from the maximally compact component $\mathcal{K}$ of the group intertwined by an element from a maximal Abelian subgroup $\mathcal{A}$: $G=KAK^{\prime},\,K,K^{\prime}\in\mathcal{K},\,A\in\mathcal{A}$. Such idea of non-compact averaging has been introduced in [36, 8] in the context of averaging quantum states over SLOCC-type operations, and generalized to averaging over arbitrary finite-dimensional matrix representations of reductive groups in the recent work [28]. Following [28], let us take arbitrary matrix representation $\pi_{\mathcal{G}}$ of the group $\mathcal{G}$ and define the averaging of a quantum state with respect to the chosen representation as follows:
@@
in which the integration is defined over the Cartan “KAK” decomposition of $\mathcal{G}$. The integration over the two copies of maximally compact components $\mathcal{K},\mathcal{K}^{\prime}$ is performed uniformly, which means that $\int_{\mathcal{K}}\operatorname{d}\!K$ represents Haar integral over $\mathcal{K}$. Integration over $\mathcal{A}$ is performed with respect to any finite normalized measure on this (in general non-compact) manifold, but there appears one crucial assumption: elements of $\mathcal{A}$ enter the integral as normalized matrices: $AA_{\textrm{n}}$=A/||A||, in order to guarantee trace nonincreasing character of the map (22).
@@
$\mathcal{M}^{\Phi}_{\mathcal{G}}(\rho)=\int\operatorname{d}\!G\,\mathcal{T}_{G}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{G}^{\operatorname{in}\dagger}(\rho),$ (24)
@@
In analogy with (2) the Choi matrix of a channel twirled with respect to a group $\mathcal{G}$ reads:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\left(\mathcal{M}^{\Phi}_{\mathcal{G}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
$=$ $\left(\int\operatorname{d}\!G(\mathcal{T}_{G}^{\operatorname{out}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\mathcal{T}_{G}^{\operatorname{in}\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\right)\rho_{\operatorname{ME}}$
$=$ $\int\operatorname{d}\!G\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left(\pi_{\mathcal{G}}^{\operatorname{in}}(G)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}\left(\pi_{\mathcal{G}}^{\operatorname{in}}(G)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\right]\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
Now all of the derivations (6)–(2) concerning the unitary case hold, since they do not utilize the property of unitarity of the underlying representations, and we have the following expression for the Choi matrix of the channel (24):
@@
In full analogy with the previous case one may want to avoid using contragradient representations in the generalized Schur-Weyl duality [^3] Note that although formally $\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}$ is not a contragradient representation to $\pi_{\mathcal{G}}^{\operatorname{in}}$, for a non-unitary representation $\pi_{\mathcal{G}}^{\operatorname{in}}$, the restriction $\bar{\pi}_{\mathcal{K}}^{\operatorname{in}}$ is contragradient to $\pi_{\mathcal{K}}^{\operatorname{in}}$, since we can without loss of generality always take unitary $\pi_{\mathcal{K}}^{\operatorname{in}}$.. Utilizing Lemma 3 the formula (27) becomes:
@@
Now the restricted representation $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$ reads in this case simply $U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, for $U\in\textrm{SU}(d)$, and we have, following (35):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)^{\Gamma},$ (38)
where now $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]$ are ordinary permutation operators corresponding to standard Schur-Weyl duality for $\textrm{U}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$ and the permutation group $\textrm{S}_{(t_{\operatorname{in}}+t_{\operatorname{out}})}$.